OpenStudy (anonymous):

use inverse operations to solve each equation. explain each step 19=h/3-8 5t+8=43

OpenStudy (lieutenantgeneral):

19=h/3-8: h = 81 because 27 - 81/3 = 8 5t+8=43: t = 7 because 5t - 35 = 0 and 5 * 7 = 35 so t = 7 = 5t

OpenStudy (anonymous):

So you have to break it down by steps. For the first one19= h over 3-8. First what's 19+8, which is 27. So if you forget the minus 8 part, the answe has to equal 27. 27 x 3 is 81. So if we plug the numbers in, 19= 81/3 -8 makes sense. For the second one, 5t+8= 43. 43- 8 is 35. What times 5 is 35, that's 7. So plug it in, 5x7 +8 = 43. Hope it makes sense now!!